### v1

# CTU-SCU/2024013

Abstract
The finiteness of the mixed state entanglement entropy $S_{\text{vN}}$ in CFT2 enables us to show that, the dynamical equation of the coincidence limit of $(\frac{1}{2}S_{vN}^{2});_{ij}$ is precisely Einstein equation. Thus the entanglement entropy induces an internal entropic gravitational geometry in CFT. The CFT RG equation turns out to be a geometric identity. In this entropic gravity, a covariant Klein-Gordon equation is found. As a natural consequence, we are led to conjecture that, the gauge/gravity duality is “ flat background QFT coupled with an internal entropic gravitational geometry on the boundary $=$ gravity and coupled QFT in the bulk”.
preprint: preprintnumbers
Introduction
The AdS/CFT correspondence [1] dominates high energy theory in last decades. As a testable realization for the holographic principle [2], it conjectures that a weakly coupled gravitational theory in the bulk of AdSd+1 is equivalent to a strongly coupled $d$ dimensional CFTd on the conformally flat boundary. Among innumerous developments, many works have been devoted to a central subject, namely reconstructing gravity from the dual QFT, or realizing Einstein equation in QFT.
Since Ryu and Takayanagi (RT) identified the pure state entanglement entropy of CFTd with the bulk minimal surface anchored on the conformal boundary of AdSd+1 [3, 4], out of many approaches, using the entanglement entropy to generate the dual geometry becomes a very promising direction. In this paper, we are working with $d=2$, in which the minimal surfaces are geodesics.
In an entangled pure CFT state $\psi_{A\bar{A}}$, entanglement entropy classifies quantum entanglement between complementary parts $A$ and $\bar{A}$ , defined by the von Neumann entropy
$S_{\text{vN}}(A)=-\mathrm{Tr}\rho_{A}\log\rho_{A},$ (1)
for a reduced density matrix $\rho_{A}=\mathrm{Tr}_{\bar{A}}|\psi_{A\bar{A}}\rangle\langle\psi_{A\bar{A}}|$. However, both the pure state entanglement entropies and the boundary anchored geodesics are infinite. This asymptotic behavior makes it very difficult, if not completely impossible, to get certain exact relations.
It is of importance to note that, for disjoint subsystems $A$ and $B$ in a mixed state $\rho_{AB}$, the entanglement entropy defined in eqn. (1) fails to characterize entanglement between $A$ and $B$. The traditional resolution is to purify the mixed state by adding auxiliary systems $\bar{A}\bar{B}$. Then the entanglement of purification (EoP or $E_{P}$) [5] is defined by minimizing the entanglement entropy $S_{\text{vN}}(A\bar{A}:B\bar{B})$ over all possible purifications, and quantifies the entanglement between $A$ and $B$: $S_{\mathrm{vN}}(A:B):=E_{P}=\min_{\bar{A}\bar{B}}S_{\text{vN}}(A\bar{A}:B\bar{B})$. The EoP is then conjectured to equals the entanglement wedge cross-section (EWCS or $E_{W}$) in the AdS bulk [6]. However, minimizing EoP over all possible purifications poses significant challenges in verifying $E_{P}=E_{W}(A:B)$.
In a recent work, we proposed an alternative to purification in CFT by subtracting the undetectable regions [7]. The entanglement entropy $S_{\mathrm{vN}}(A:B)$ now can be easily calculated and we do confirm the exact equality
$S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right).$ (2)
The most important property is that both quantities are finite, and the bulk geodesics have not to be attached on the infinitely far away conformal boundary. This is generic, since in multipartite systems, for any two disjoint parts $A$ and $B$, we always have finite $S_{\mathrm{vN}}\left(A:B\right)$ and $E_{W}\left(A:B\right)$. When combined with the Synge’s world function method, in this paper, we are able to derive the dynamical equation for entanglement entropy, which turns out to be precisely the Einstein equation.
The dynamical equation of entanglement entropies
Given a metric, we know how to compute the geodesic distance between two points. On the other hand, the metric can also be extracted from geodesics. Consider a geodesic $x(\tau),\tau\in[0,t]$ connecting two points $x(\tau=0)$ and $x^{\prime}(\tau=t)$. $L(x,x^{\prime})$ is the geodesic length. The Synge’s world function [8] is a biscalar of $x$ and $x^{\prime}$, defined as half the squared distance,
[…]
Denote the coincidence limit
$[\Omega]:=\lim_{x^{\prime}\to x}\Omega(x,x^{\prime}),$ (4)
for any bitensor $\Omega(x,x^{\prime})$. The metric turns out to be
$g_{\mu\nu}=-[\sigma_{\mu\nu^{\prime}}]:=-\lim_{x^{\prime}\to x}\partial_{{x^{\prime}}^{\nu}}\partial_{{x}^{\mu}}\sigma(x,x^{\prime}),$ (5)
or
$g_{\mu\nu}=[\sigma_{\mu\nu}]:=\lim_{x^{\prime}\to x}\nabla_{x^{\mu}}\partial_{{x}^{\nu}}\sigma(x,x^{\prime}).$ (6)
Note $\sigma_{\mu}:=\partial_{x^{\mu}}\sigma$ is a vector with respect to point $x$, but is a scalar for point $x^{\prime}$. So, $\sigma_{\mu\nu}:=\nabla_{\nu}\partial_{\mu}\sigma$ involves covariant derivatives, whereas $\sigma_{\mu\nu^{\prime}}:=\partial_{\nu^{\prime}}\partial_{\mu}\sigma$ does not. We thus conclude
$\mathrm{Metric}\Leftrightarrow\mathrm{Geodesic}$ (7)
are equivalent.
With respect to metric, we have Einstein action and equation
$S_{E}=\frac{1}{16\pi G}\int d^{D}x\,\sqrt{-g}(R-2\Lambda),$ (8)
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}\,R+\Lambda g_{\mu\nu}=0,$ (9)
where $D=d+1$ is the spacetime dimension. We only address vacuum with $\Lambda\not=0$. Thus contracting $g^{\mu\nu}$ with eqn. (9), we find $R=\frac{2D}{D-2}\Lambda$ and the Einstein equation can also be put into
$R_{\mu\nu}-\frac{1}{D}g_{\mu\nu}\,R=0.$ (10)
EOM for the geodesics
We are now ready to write down the equation for the geodesics. Since
$[\sigma_{\alpha\mu\beta\nu}]=\frac{1}{3}\Big{(}R_{\alpha\nu\beta\mu}+R_{\alpha\beta\nu\mu}\Big{)},$ (11)
one finds
$\Sigma_{\mu\nu}:=[\sigma^{\alpha\beta}\,\sigma_{\alpha\mu\beta\nu}]=g^{\alpha\beta}[\sigma_{\alpha\mu\beta\nu}]=\frac{1}{3}R_{\mu\nu},$ (12)
and
$\Sigma:=[\sigma^{\mu\nu}\sigma^{\alpha\beta}\,\sigma_{\alpha\mu\beta\nu}]=g^{\mu\nu}\Sigma_{\mu\nu}=\frac{1}{3}R.$ (13)
Note raising or lowering indices and taking the limit commute. The geodesic version of Einstein equation is
$\Sigma_{\mu\nu}-\frac{1}{D}[\sigma_{\mu\nu}]\,\Sigma=0,$ (14)
taking the identical form as the metric Einstein equation (10).
The dynamics of Entanglement Entropy
To study the mixed state entanglement entropy in CFT2, referring to Fig. (1), we take a time slice of an infinite bipartite system. Segments $A,B$ are separated by regions $C,D$. The Brown-Henneaux’s formula [9] $c=3R_{\text{AdS}}/2G^{(3)}$ is assumed. $\gamma_{1},\gamma_{2}$ and $L_{AB}$ are all geodesics. RT formula gives $\gamma_{1}=S_{vN}(C:ABD)$ and $\gamma_{2}=S_{vN}(D:ABC)$, where we also use the geodesic itself to denote its length. From the Ultra-parallel theorem [10] (Theorem 1.1.6)) in hyperbolic geometry, the shortest curve connects $\gamma_{1}$ and $\gamma_{2}$ is a unique geodesic $L_{AB}$. This geodesic is identified as EWCS, $E_{W}(A:B):=L_{AB}$. From the figure, we see $\xi_{1}$ and $\xi_{2}$ fix $\gamma_{1}$, $\xi_{3}$ and $\xi_{4}$ fix $\gamma_{2}$, $L_{AB}$ is completely fixed by $\gamma_{1}$ and $\gamma_{2}$. Therefore, $(x(\xi_{i}),z(\xi_{i}))$ and $(x^{\prime}(\xi_{i}),z^{\prime}(\xi_{i}))$ are functions of $\xi_{i}$’s. Obviously $E_{W}(A:B)$ is a finite quantity and the end points of $L_{AB}$ move freely in the bulk accompanied with the variations of $A,B$ accordingly. The picture also demonstrates explicitly how an extra dimension arises from the boundary theory: four points on the boundary leads to two points in the bulk. It is not hard to imagine the same pattern holds also for higher dimensional theories.
Figure 1: Bipartite mixed state in CFT2 on a time slice. Geodesic $\gamma_{1}$ is completely fixed by $\xi_{1}$ and $\xi_{2}$, geodesic $\gamma_{2}$ is completely fixed by $\xi_{3}$ and $\xi_{4}$. From the Ultra-parallel theorem in hyperbolic geometry, as the shortest curve connecting $\gamma_{1}$ and $\gamma_{2}$, $L_{AB}$ is a geodesic and unique. $(x,z)$ and $(x^{\prime},z^{\prime})$ are completely fixed by all four $\xi_{i}$’s. The EWCS of $A,B$ is defined as $L_{AB}$, which equals the CFT entanglement entropy between segments $A$ and $B$, i.e. $S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right)=L_{AB}$.
As proposed in [6], and proved in [7], the entanglement between $A$ and $B$ exactly matches the EWCS
$S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right)=L_{AB}=\frac{c}{6}\,\log\Big{(}1+2\alpha+2\sqrt{\alpha(\alpha+1)}\Big{)},$ (15)
with the cross ratio
$\alpha=\frac{(\xi_{3}-\xi_{2})(\xi_{4}-\xi_{1})}{(\xi_{2}-\xi 1)(\xi_{4}-\xi_{3})}.$ (16)
It is easy to see that at the limit $C,D\to 0$, the mixed state finite quantities reduce to the well known pure state infinite one: $\frac{c}{3}\log\frac{\ell}{\epsilon}$, where $\ell$ is the length of the segment $A$ and $\epsilon$ is a UV cutoff.
We now have all the ingredients to achieve our main purpose. From eqns. (5) and (15), it becomes trivial to get the metric of the dual bulk geometry from the CFT2 mixed state entanglement entropy $S_{\mathrm{vN}}(A:B)$, which of course is AdS3. Using eqn. (5) but not (6) because to get the metric, covariant derivatives should be avoided. This had been a highly non-trivial problem when use the pure state entanglement entropies [11, 12, 13].
The most important thing is that, from eqns. (14) and (15), we now have a dynamical equation for the entanglement entropy! Defining
$\chi:=\frac{1}{2}S_{\mathrm{vN}}^{2},\quad g_{ij}:=[\chi_{ij}],\quad R_{ij}:=\left[\chi^{k\ell}\chi_{ki\ell j}\right],\quad R:=[\chi^{ij}R_{ij}],$ (17)
where the derivatives $i,j\cdots$ are with respect to four $\xi_{i}$’s, we are ready to obtain the dynamic equation of CFT entanglement entropy
$R_{ij}-\frac{1}{D}g_{ij}R=0,$ (18)
which is precisely the Einstein equation! Since all the quantities and parameters in this equation belong to CFT, it is indeed a CFT dynamical equation for the entanglement entropy.
It is also easy to write down the action of the entanglement entropy
$S=\frac{3}{16\pi G}\int\,d^{D}x_{i}\,\sqrt{g_{ij}}\,(R-\frac{2}{3}\Lambda)=\frac{3}{16\pi G}\Big{[}\int\,d^{D}x_{i}\,\sqrt{\chi_{ij}}\,(\chi^{k\ell}\chi^{mn}\chi_{km\ell n}-\frac{2}{3}\Lambda)\Big{]}.$ (19)
It is intriguing that at least for $D=3$, the coupling $3/16\pi G$ becomes $\frac{3c}{8\pi R_{AdS}}$, which is a strong coupled theory for large central charge $c$.
Before moving on, we should make some comments about on which $\xi_{i}$’s the coincidence limit and derivatives act. Referring to Fig.(1), for the bulk geodesic equations, all indices $\alpha,\beta,\cdots$ in eqns (11-14) take values in $(x,z)$, and the coincident limit is $(x^{\prime}\to x,z^{\prime}\to z)$. Whereas for the CFT entanglement entropy, since all the equations are tensorial, coordinate transformations enable the derivatives in eqns (17-19) are with respect to $\xi_{i}$’s directly. Without loss of generality, we can choose $i,j\cdots$ taking values in ($\xi_{1}$,$\xi_{2}$), then accordingly, taking the coincidence limit is to set $\big{\{}x_{4}\to x_{1}$ and $x_{3}\to x_{2}$ $(A,B\to 0)\big{\}}$.
Note though we find the Einstein equation for entanglement entropy from the mixed bipartite system, the derivation is completely generic and works for any multipartite mixed state. The key point is that we have finite entanglement entropies whose end points are free to move in the bulk.
Interpretation, RG equation and Klein-Gordon equation
Expand Gauge/Gravity duality
We now see that, for the Gauge/Gravity, on the QFT side, we have a hidden entropic geometry in addition to the CFT itself. Whereas on the gravity side, we have spacetime geometry and QFT living on it. The spacetime geometry and entropic geometry have identical structures. It seems that our previous understanding of Gauge/Gravity duality is incomplete. Either side has both geometry and QFT.
Another point is that the internal entropic geometry could be interpreted as generated by the entanglement entropy. The CFT and the internal entropic geometry are linked by the entanglement entropy. Once a QFT quantity is expressed in terms of entanglement entropy, one can use the geometry results to study it. Moreover, thermodynamic property of gravity becomes a natural consequence.
Basically, there are three fundamental equations, Einstein equation, covariant Klein-Gordon equation, RG equation. If the above expansion of Gauge/Gravity duality is correct, all three equations should appear in both sides and correspond one-to-one. We already know that in bulk gravity, Einstein equation and covariant Klein-Gordon equation exist, whereas in the CFT on the boundary, we have entropic Einstein equation and RG equation. Therefore, we need to find the covariant Klein-Gordon equation in CFT and RG equation in gravity.
RG equation as a geometric identity
For a classically scale-invariant theory living on a $D$-dimensional manifold $\mathcal{M}$ with the metric $ds^{2}=\gamma_{ab}dx^{a}dx^{b}$, the Callan-Symanzik RG equation is [14]:
$\left[\ell\frac{\partial}{\partial\ell}-2\int_{\mathcal{M}}\gamma^{ab}\frac{\delta}{\delta\gamma^{ab}}\right]\log Z_{\text{CFT}}=0,$ (20)
where $\ell$ is the length scale. Using the replica trick [15] and Rényi entropy
$S_{\text{vN}}^{(n)}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{M}_{n}}}{\left(Z_{\mathcal{M}}\right)^{n}}\right],$ (21)
where $\mathcal{M}_{n}$ is the replicated manifold, one can show the RG equation for the entanglement entropy is [16]
$\ell\frac{\partial}{\partial\ell}S_{\text{vN}}=\frac{c}{6}.$ (22)
The constant $\frac{c}{6}$ on the right hand side is for $D=3$.
On the other hand, in any geometry, there is an identity for the world function (3),
$g^{\mu\nu}\sigma_{\mu}\sigma_{\nu}=2\sigma,\quad\sigma(x,x^{\prime})=\frac{1}{2}L^{2}(x,x^{\prime})$ (23)
In AdS3, we still consider a time slice whose isometry is SL$(2,R)$. Those isometries preserve the distance between points. Except the vertical ones, any other geodesic is a segment of a semi-circle,
$(x-x_{0})^{2}+z^{2}=r^{2}.$ (24)
We can map a semi-circle to a vertical line by the following isometries
$(x,z)$$\rightarrow$ $(x-x_{0}+r,z)\rightarrow(\frac{x-x_{0}+r}{(x-x_{0}+r)^{2}+z^{2}},\frac{z}{(x-x_{0}+r)^{2}+z^{2}})$ (25)
$\rightarrow$ $(\tilde{x},\tilde{z}):=\Big{(}\frac{x-x_{0}+r}{(x-x_{0}+r)^{2}+z^{2}}-\frac{1}{2r},\frac{z}{(x-x_{0}+r)^{2}+z^{2}}\Big{)}.$ (26)
In terms of $(\tilde{x},\tilde{z})$, the semi-circle is a vertical line ${\sigma}={\sigma}(0,\tilde{z}_{1};0,\tilde{z}_{2})$. Now, apply the geometric identity (23),
$z^{2}\left[(\partial_{x}\sigma)^{2}+(\partial_{z}\sigma)^{2}\right]=\tilde{z}^{2}(\partial_{\tilde{z}}{\sigma})^{2}=2{\sigma}.$ (27)
Obviously, $\tilde{z}$ is the renormalization scale $\ell$. Including the coupling $c=3R_{\text{AdS}}/2G^{(3)}$, in either the bulk spacetime gravity or CFT internal entropic geometry, $\sigma=\frac{1}{2}L^{2}=\frac{1}{2}S_{\text{vN}}^{2}$, we thus have in both bulk geometry and boundary CFT
$\ell\frac{\partial}{\partial\ell}S_{\text{vN}}=\frac{c}{6},$ (28)
precisely the RG equation (22).
Covariant Klein-Gordon equation in CFT
Usually the entanglement entropy is computed by using the replica method, and the Rényi entropy is determined by the partition function of the replicated manifold. It is generally believed that the replicated partition function equals the correlation function of twisted operators,
$G\left(u(\xi_{i}),v(\xi_{i})\right)=\mathrm{Tr}_{A}\rho_{A}^{n}.$ (29)
In a recent work [7], we have obtained that
$\mathrm{Tr}_{A}\rho_{A}^{n}=e^{\frac{1}{2}(\frac{1}{n}-n)S_{\mathrm{vN}}\left(A:B\right)}.$ (30)
So,
$G\left(u(\xi_{i}),v(\xi_{i})\right)=e^{\frac{1}{2}(\frac{1}{n}-n)S_{\mathrm{vN}}\left(A:B\right)},$ (31)
Note $u\left(\xi_{i}\right)$ and $v\left(\xi_{i}\right)$ are basically $(x,z)$ and $(x^{\prime},z^{\prime})$ respectively in Fig. (1), but located on the boundary of the cut geometry. Thus $G\left(u(\xi_{i}),v(\xi_{i})\right)$ is not the usual CFT propagator but effectively a four point function.
Therefore, from the entropy Einstein equation (18) for $S_{\mathrm{vN}}$, including the temporal direction, one can show that
$(\nabla^{2}-m^{2})G\left(u(\xi_{i}),v(\xi_{i})\right)=\delta^{3}_{\cal M}(\xi_{i}-\xi_{j}),$ (32)
where $\triangle_{n}$ is the weight of the twist field and mass $m=\sqrt{\triangle_{n}\left(\triangle_{n}-2\right)}\overset{\triangle_{n}\to\infty}{=}\triangle_{n}$ at large $c$ limit. One should be careful about grouping $\xi_{i}$’s as we mentioned in the last section. For example, we can set $\xi_{i}\in(\xi_{1},\xi_{2})$ and $\xi_{j}\in(\xi_{3},\xi_{4})$. This covariant equation is a manifestation of the entropic geometry. Note all quantities in the CFT covariant propagator (31) are expressed in terms of CFT parameters. The usual CFT propagator is the limit of this covariant propagator.
Summary and discussion
By using RT formula to identify geodesics with entanglement entropies, we showed that the dynamics of entanglement entropy in CFT2 are precisely Einstein equation. We thus uncovered a hidden geometry in CFT. We confirmed that in addition to Einstein equation and covariant Klein-Gordon equation, in the bulk, a geometrical identity is the RG equation. On the CFT side, in addition to the entropic Einstein equation, we also gave the covariant Klein-Gordon equation, showed that the RG equation is an identity of the entropic geometry.
All those observations lead us to believe that we should expand the interpretation of Gauge/Gravity duality to “flat QFT + internal entropic geometry $=$ spacetime geometry + QFT in curved background”. In this understanding, ER$=$EPR becomes manifest. Gravity as entropic force is also apparent.
Since the ordinary QFT and entropic gravity are linked through entanglement entropy, we could study what the structures of the entropic geometry, such as black holes, wormholes $\cdots$, are in QFT without much difficulty.
In physics, energy dominates and links everything. But now, we tend to believe the fundamental object in physics is entropy.
Acknowledgements.
Acknowledgments.
This work is supported in part by NSFC (Grant No. 12105191, 12275183 and 12275184).