### v4

# CTP-SCU/2024013

Abstract
The finiteness of the mixed state entanglement entropies enables us to show that, the dynamical equation of the entanglement entropy in CFT2 is precisely three dimensional Einstein’s equation. A profound relation between the cosmological constant and CFT2 entanglement entropy is given. Thus entanglement entropies induce internal gravitational geometries in CFT2. Deriving the dual metric from an entanglement entropy becomes a trivial procedure. Surprisingly, we find that the renormalization group equation is a geometric identity.
preprint: preprintnumbers
Introduction
The AdS/CFT correspondence [1] dominates high energy theory in the last three decades. As a testable realization for the holographic principle [2], it conjectures that a weakly coupled gravitational theory in the bulk of AdSd+1 is equivalent to a strongly coupled CFTd on the conformally flat boundary. Among numerous developments, a central subject is to reconstruct gravity from the dual QFT, or “derive” Einstein’s equation in QFT.
Since Ryu and Takayanagi (RT) identified the pure state entanglement entropy of CFTd with the area of bulk minimal surfaces anchored on the conformal boundary of AdSd+1 [3, 4], among many approaches, using entanglement entropy to generate the dual geometry has become a very promising direction. For $d=2$ considered in this letter, the minimal surfaces are geodesics.
In an entangled pure CFT state $\psi_{A\bar{A}}$, entanglement entropy classifies quantum entanglement between complementary parts $A$ and $\bar{A}$. It is defined by the von Neumann entropy: $S_{\text{vN}}(A)=-\mathrm{Tr}\rho_{A}\log\rho_{A}$. Here $\rho_{A}$ is a reduced density matrix $\rho_{A}=\mathrm{Tr}_{\bar{A}}|\psi_{A\bar{A}}\rangle\langle\psi_{A\bar{A}}|$. However, both the pure state entanglement entropies and lengths of the boundary-anchored geodesics are divergent. This asymptotic behavior makes it very difficult, if not impossible, to extract certain exact relations.
It is crucial to note that, the entanglement entropy defined above fails to characterize the entanglement in a mixed state $\rho_{AB}$. The traditional resolution is to purify the mixed state by adding auxiliary systems $\bar{A}\bar{B}$. Then the entanglement of purification (EoP or $E_{P}$) [5] is defined by minimizing $S_{\text{vN}}(A\bar{A}:B\bar{B})$ over all possible purifications, and quantifies the entanglement between $A$ and $B$: $S_{\mathrm{vN}}(A:B):=E_{P}=\min_{\bar{A}\bar{B}}S_{\text{vN}}(A\bar{A}:B\bar{B})$. As an application of RT formula on mixed states, the EoP is conjectured to equals the entanglement wedge cross-section (EWCS or $E_{W}$) in the AdS bulk [6]. However, optimization over purifications is, in practice, not workable in CFTs.
Recently, the SUBTRACTION approach is proposed to calculate the mixed state entanglement entropy $S_{\mathrm{vN}}\left(A:B\right)$ in CFT [7, 8, 9]. It proves that $S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right)$ is an exact equation. The most important property is that both quantities are finite for disjoint subsystems $A$ and $B$, and the bulk geodesics of $E_{W}\left(A:B\right)$ do not have to be attached on the conformal boundary. This enables us to explicitly show how three dimensional Einstein’s equation emerges from CFT2 in this paper.
The line of our reasoning is as follows:
  1. 1.
    Metric $\Leftrightarrow$ Geodesic. They can be derived from each other.
  2. 2.
    (Metric version) Einstein’s equation $\Rightarrow$ Geodesic version Einstein’s equation.
  3. 3.
    RT formula: (Mixed state) Entanglement entropy $=$ Geodesic length.
    Since both quantities are finite and this equation is exact [7, 8, 9], we have:
  4. 4.
    Geodesic version Einstein’s equation $\Rightarrow$ Entanglement Entropy version Einstein’s equation.
Then we will show that the renormalization group (RG) equation is surprisingly a geometric identity.
Einstein’s equation from entanglement entropies in CFT2
Step one: Metric $\Leftrightarrow$ Geodesic
Given a metric, we know how to compute the geodesic distance between two points. On the other hand, the metric can also be extracted from geodesics. Consider a geodesic $x(\tau),\tau\in[0,t]$ connecting two points $x(\tau=0)$ and $x^{\prime}(\tau=t)$. $L(x,x^{\prime})$ is the geodesic length. The Synge’s world function is a biscalar of $x$ and $x^{\prime}$, defined as,
[…]
The metric turns out to be [10]
$g_{\mu\nu}=-[\sigma_{\mu\nu^{\prime}}]:=-\lim_{x^{\prime}\to x}\partial_{{x^{\prime}}^{\nu}}\partial_{{x}^{\mu}}\sigma(x,x^{\prime}),$ (2)
or
$g_{\mu\nu}=[\sigma_{\mu\nu}]:=\lim_{x^{\prime}\to x}\nabla_{x^{\mu}}\partial_{{x}^{\nu}}\sigma(x,x^{\prime}),$ (3)
where []delimited-[][$\cdots$][ ⋯ ] denote the coincidence limit $x\to x^{\prime}$. By default, the indices of $\sigma$ are (covariant) derivatives. Note $\sigma_{\mu}:=\partial_{x^{\mu}}\sigma$ is a vector with respect to point $x$, but is a scalar for point $x^{\prime}$. So, $\sigma_{\mu\nu}:=\nabla_{\nu}\partial_{\mu}\sigma$ involves covariant derivatives, whereas $\sigma_{\mu\nu^{\prime}}:=\partial_{\nu^{\prime}}\partial_{\mu}\sigma$ does not. We thus conclude
$\mathrm{Metric}\Leftrightarrow\mathrm{Geodesic}$ (4)
Step two: Geodesic version Einstein’s equation
With respect to metric, we have Einstein’s action and equation
$S_{E}=\frac{1}{16\pi G}\int d^{d+1}x\,\sqrt{-g}(R-2\Lambda),$ (5)
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}\,R+\Lambda g_{\mu\nu}=0.$ (6)
We only address vacuum with $\Lambda\not=0$. The matter sourced theories can be obtained by straightforward generalization. Based on step one, we are ready to write down the equation for the geodesics. Since
$[\sigma_{\alpha\mu\beta\nu}]:=\lim_{x^{\prime}\to x}\nabla_{\nu}\nabla_{\beta}\nabla_{\mu}\partial_{\alpha}\sigma=\frac{R_{\alpha\nu\beta\mu}+R_{\alpha\beta\nu\mu}}{3},$ (7)
one finds
$\Sigma_{\mu\nu}:=[\sigma^{\alpha\beta}\,\sigma_{\alpha\mu\beta\nu}]=g^{\alpha\beta}[\sigma_{\alpha\mu\beta\nu}]=\frac{1}{3}R_{\mu\nu},$ (8)
and
$\Sigma:=[\sigma^{\mu\nu}\sigma^{\alpha\beta}\,\sigma_{\alpha\mu\beta\nu}]=g^{\mu\nu}\Sigma_{\mu\nu}=\frac{1}{3}R.$ (9)
Note raising or lowering indices and taking the coincidence limit commute. The geodesic version of Einstein’s equation is
$\Sigma_{\mu\nu}-\frac{1}{2}[\sigma_{\mu\nu}]\,\Sigma+\frac{\Lambda}{3}[\sigma_{\mu\nu}]=0,$ (10)
taking an identical form as Einstein’s equation (6).
Step three: RT formula, Mixed state EE $=$ Geodesic
In [6], it is proposed that the mixed state entanglement entropy $S_{\mathrm{vN}}\left(A:B\right)$ for disjoint regions $A$ and $B$ in CFT equals the EWCS ($E_{W}\left(A:B\right)$) in the bulk. Since the EWCS of disjoint regions is finite, this equation should be exact. Using the subtraction approach, we calculated the mixed state entanglement entropy $S_{\mathrm{vN}}\left(A:B\right)$ in [7] for static case, in [8] for time-dependent covariant case and in [9] for thermal CFT2. The exactness of $S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right)$ has been confirmed.
For simplicity and without loss of generality, referring to Figure 1, we take a time slice of an infinite system as an illustration. This is the scenario addressed in [7]. Including time dependence or extending to other CFT2 can be easily done by using the results in [8, 9]. Segments $A,B$ are separated by regions $C,D$. The Brown-Henneaux’s formula [11] $c=3R_{\text{AdS}}/2G^{(3)}$ is assumed. $\gamma_{1},\gamma_{2}$ and $L_{AB}$ are all geodesics. RT formula gives $\gamma_{1}=S_{\text{vN}}(C:ABD)$ and $\gamma_{2}=S_{\text{vN}}(D:ABC)$, where we also use the geodesic itself to denote its length. From the Ultra-parallel theorem [12] (Theorem 1.1.6)) in hyperbolic geometry, the shortest curve connects $\gamma_{1}$ and $\gamma_{2}$ is a unique geodesic $L_{AB}$. This geodesic is identified as EWCS, $E_{W}(A:B):=L_{AB}$. From the figure, we see $\xi_{1}$ and $\xi_{2}$ fix $\gamma_{1}$, $\xi_{3}$ and $\xi_{4}$ fix $\gamma_{2}$, $L_{AB}$ is completely fixed by $\gamma_{1}$ and $\gamma_{2}$. Therefore, $(x(\xi_{i}),z(\xi_{i}))$ and $(x^{\prime}(\xi_{i}),z^{\prime}(\xi_{i}))$ are functions of $\xi_{i}$’s. Obviously $E_{W}(A:B)$ is a finite quantity and the end points of $L_{AB}$ move freely in the bulk accompanied with the variations of $A,B$ accordingly. The picture also demonstrates explicitly how an extra dimension arises from the boundary theory: four points on the boundary leads to two points in the bulk. For the time-dependent covariant scenario, four points on the boundary have eight parameters. However, two of these turn out to be gauge parameters. The remaining six independent parameters exactly match two points in the 3-dimensional bulk [8].
Figure 1: Bipartite mixed state in CFT2 on a time slice. Geodesic $\gamma_{1}$ is completely fixed by $\xi_{1}$ and $\xi_{2}$, geodesic $\gamma_{2}$ is completely fixed by $\xi_{3}$ and $\xi_{4}$. $L_{AB}$ is a geodesic and unique. $(x,z)$ and $(x^{\prime},z^{\prime})$ are completely fixed by all four $\xi_{i}$’s. The EWCS of $A,B$ is defined as $L_{AB}$, which equals the CFT entanglement entropy between segments $A$ and $B$, i.e. $S_{\mathrm{vN}}\left(A:B\right)=E_{W}\left(A:B\right)=L_{AB}$.
So, in AdS3/CFT2, for a mixed state $\rho_{AB}$, we have an exact equation
$S_{\mathrm{vN}}\left(A:B\right)=\frac{c}{6}\log\Big{|}1+2\alpha+2\sqrt{\alpha(\alpha+1)}\Big{|}=E_{W}\left(A:B\right),$ (11)
where the cross ratio $\alpha$ is determined by the specific CFT. For example,
$\alpha=\frac{(\xi_{3}-\xi_{2})(\xi_{4}-\xi_{1})}{(\xi_{2}-\xi_{1})(\xi_{4}-\xi_{3})},$ (12)
for the case illustrated in Figure 1. At the limit $C,D\to 0$, the mixed state finite quantities reduce to the well known pure state infinite one: $\frac{c}{3}\log\frac{\ell}{\epsilon}$, where $\ell$ is the length of the segment $A$ and $\epsilon$ is a UV cutoff.
We emphasize that, although we consider only the infinite bipartite system here, the derivations are completely generic and work for any multipartite, excited mixed states. The key point is that we have finite mixed state entanglement entropies and the dual geodesics are not restricted on the boundary.
Step four: Einstein’s equation from CFT2
We now have all the ingredients to achieve our main purpose. From eqns. (2) and (11), it becomes trivial to extract the metric of the dual bulk geometry from the CFT2 mixed state entanglement entropy $S_{\mathrm{vN}}(A:B)$. Using eqn. (2) but not (3) because covariant derivatives cannot be defined prior to the metric. This had been a highly non-trivial problem when using the divergent pure state entanglement entropies [13, 14, 15].
The most important thing is that, from eqns. (10) and (11), we now have a dynamical equation for CFT2 entanglement entropies! Defining
$\chi$ ::=:= $\frac{1}{2}S_{\mathrm{vN}}^{2},\quad\quad g_{ij}:=[\chi_{ij}],$
$R_{ij}$ ::=:= $\left[\chi^{k\ell}\chi_{ki\ell j}\right],\quad R:=[\chi^{ij}R_{ij}],$ (13)
where the derivatives $i,j\cdots$ are with respect to four $\xi_{i}$’s, we immediately obtain
$R_{ij}-\frac{1}{2}g_{ij}R+\frac{\Lambda}{3}g_{ij}=0,$ (14)
which is precisely Einstein’s equation! Since all the quantities and parameters in this equation belong to CFT2, it is truly a CFT dynamical equation for the entanglement entropy. Moreover, after contracting with $g^{ij}$, a profound relation between the cosmological constant and CFT entropy is found
[…]
So, the cosmological constant is not really a free parameter, but determined by the CFT entanglement entropy. The action of the entanglement entropy is
$I$ $=$ $\frac{3}{16\pi G}\int\,d^{d+1}x_{i}\,\sqrt{g_{ij}}\,(R-\frac{2}{3}\Lambda)$
$=$ $\frac{3}{16\pi G}\Big{[}\int\,d^{d+1}x_{i}\,\sqrt{\chi_{ij}}\,(\chi^{k\ell}\chi^{mn}\chi_{km\ell n}-\frac{2}{3}\Lambda)\Big{]}.$
From eqn. (13) and this action, we see that the CFT2 induced metric is a derived but not a fundamental quantity. Especially the metric is the second derivative of entanglement entropy. This perhaps is the reason why gravity is non-renormalizable.
We should make some comments about on which $\xi_{i}$’s the coincidence limit and derivatives act. Referring to Figure (1), for the bulk geodesic equations, all greek indices $\alpha,\beta,\cdots$ in eqns (7-10) take values in $(x,z)$, and the coincident limit is $(x^{\prime}\to x,z^{\prime}\to z)$. Whereas for the CFT entanglement entropy, since all the equations are tensorial, coordinate transformations enable the derivatives in eqns (13-Einstein’s equation from entanglement entropies in CFT2) to be with respect to $\xi_{i}$’s directly. Without loss of generality, one can choose $i,j\cdots$ taking values in ($\xi_{1}$, $\xi_{2}$), then accordingly, taking the coincidence limit is to set $\big{\{}x_{4}\to x_{1}$ and $x_{3}\to x_{2}$ $(A,B\to 0)\big{\}}$. In a companion work [16], we give an explicit example of how to group the boundary indices $\xi_{i}$’s.
RG equation and Klein-Gordon equation
On the QFT side, we have an entanglement generated gravity in addition to the CFT itself. Whereas on the gravity side, we have spacetime geometry and QFT living on it. The spacetime geometry and entanglement generated gravity have identical structures.
Basically, there are three fundamental equations, Einstein’s equation, covariant Klein-Gordon equation and RG equation. We already know that in bulk gravity, Einstein’s equation and covariant Klein-Gordon equation exist, whereas in the CFT on the boundary, we have quantum entanglement generated Einstein’s equation and RG equation. Therefore, we want to find the covariant Klein-Gordon equation in CFT and RG equation in gravity.
RG equation as a geometric identity
For a classically scale-invariant theory living on a $d+1$-dimensional manifold $\mathcal{M}$ with the metric $ds^{2}=\gamma_{ab}dx^{a}dx^{b}$, the Callan-Symanzik RG equation is [17]:
$\left[\ell\frac{\partial}{\partial\ell}-2\int_{\mathcal{M}}\gamma^{ab}\frac{\delta}{\delta\gamma^{ab}}\right]\log Z_{\text{CFT}}=0,$ (17)
where $\ell$ is the length scale. Using the replica trick [18] and Rényi entropy
$S_{\text{vN}}^{(n)}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{M}_{n}}}{\left(Z_{\mathcal{M}}\right)^{n}}\right],$ (18)
where $\mathcal{M}_{n}$ is the replicated manifold, one can show the RG equation for the entanglement entropy is [19]
$\ell\frac{\partial}{\partial\ell}S_{\text{vN}}=\frac{c}{6}.$ (19)
The constant $\frac{c}{6}$ on the right hand side is for $d=2$.
On the other hand, in any geometry, not limited to gravity, there is an identity for the world function (1),
$g^{\mu\nu}\sigma_{\mu}\sigma_{\nu}=2\sigma,\quad\sigma(x,x^{\prime})=\frac{1}{2}L^{2}(x,x^{\prime}).$ (20)
On a time slice of AdS3, the isometry is SL$(2,R)$. Those isometries preserve the distance between points. Except the vertical ones, any other geodesic is a segment of a semi-circle,
$(x-x_{0})^{2}+z^{2}=r^{2}.$ (21)
We can map a semi-circle to a vertical line by the following isometries
$(\tilde{x},\tilde{z}):=(\frac{x-x_{0}+r}{(x-x_{0}+r)^{2}+z^{2}}-\frac{1}{2r},\frac{z}{(x-x_{0}+r)^{2}+z^{2}}).$
In terms of $(\tilde{x},\tilde{z})$, the semi-circle is a vertical line ${\sigma}={\sigma}(0,\tilde{z}_{1};0,\tilde{z}_{2})$. Now, apply the geometric identity (20),
$z^{2}\left[(\partial_{x}\sigma)^{2}+(\partial_{z}\sigma)^{2}\right]=\tilde{z}^{2}(\partial_{\tilde{z}}{\sigma})^{2}=2{\sigma}.$ (22)
Obviously, $\tilde{z}$ is the renormalization scale $\ell$. Including the coupling $c=3R_{\text{AdS}}/2G^{(3)}$, in either the quantum entanglement generated gravity in CFT2 or the bulk spacetime gravity, $\sigma=\frac{1}{2}L^{2}=\frac{1}{2}S_{\text{vN}}^{2}$, we thus have in both boundary CFT2 and bulk geometry,
$\ell\frac{\partial}{\partial\ell}S_{\text{vN}}=\frac{c}{6},$ (23)
precisely the RG equation (19).
Covariant Klein-Gordon equation in CFT
Usually the entanglement entropy is computed by using the replica method. The Rényi entropy is determined by the partition function of the replicated manifold. The replicated partition function equals the correlation function of twisted operators,
$G\left(u(\xi_{i}),v(\xi_{i})\right)=\mathrm{Tr}_{A}\rho_{A}^{n}.$ (24)
In [7], we have obtained $\mathrm{Tr}_{A}\rho_{A}^{n}=e^{\frac{1}{2}(\frac{1}{n}-n)S_{\mathrm{vN}}\left(A:B\right)}$. So,
$G\left(u(\xi_{i}),v(\xi_{i})\right)=e^{\frac{1}{2}(\frac{1}{n}-n)S_{\mathrm{vN}}\left(A:B\right)}.$ (25)
Note $u\left(\xi_{i}\right)$ and $v\left(\xi_{i}\right)$ are basically $(x,z)$ and $(x^{\prime},z^{\prime})$ respectively in Fig. (1), but located on the boundary of the cut geometry. Thus $G\left(u(\xi_{i}),v(\xi_{i})\right)$ is not the usual CFT propagator but effectively a four point function.
To obtain the EOM of this propagator, we refer to the AdS3 bulk propagator of a scalar with mass $m^{2}=\triangle\left(\triangle-2\right)$ and $\triangle=\frac{c}{12}(n-1/n)$,
$G_{\rm Bulk}\left(u,v\right)=\frac{e^{-\triangle L\left(u,v\right)}}{1-e^{-2L\left(u,v\right)}},$ (26)
[…]
$(\nabla^{2}-m^{2})G\left(u(\xi_{i}),v(\xi_{i})\right)=\delta^{3}_{\cal M}(\xi_{i}-\xi_{j}),$ (27)
at large $c$ limit.
Summary and discussion
By using RT formula to identify geodesic lengths with entanglement entropies, we showed that the dynamics of entanglement entropy in CFT2 are precisely three dimensional Einstein equation. A profound relation between the cosmological constant and CFT entropy is also given. We thus uncovered a hidden gravitational structure in CFT. We confirmed that in addition to Einstein’s equation and covariant Klein-Gordon equation, in the bulk, a geometrical identity is the RG equation. On the CFT side, in addition to the quantum entanglement generated Einstein’s equation, we gave the covariant Klein-Gordon equation, and showed that the RG equation is a geometric identity of the emerged gravity.
Since the ordinary QFT and the induced gravity are linked through entanglement entropy, it is now possible to unify the black hole thermodynamics and Einstein’s equation.
In this paper, we have focused on AdS. Nevertheless, as long as RT formula holds, such as in dS, the derivations apply with the same fashion. This is why we did not specify the value of $\Lambda$.
Our derivations are for AdS3/CFT2. For higher dimensions, there could be two possibilities. The first one is that still the entanglement entropy satisfies Einstein’s equation. However, since the entangling surfaces are no longer points but co-dimension two surfaces, it is not obvious how to impose derivatives and then extract Einstein’s equation. The second choice, which is more promising, is to identify the dual CFT quantity of the bulk geodesics. This dual quantity automatically satisfies Einstein’s equation. It seems the story proceeds as: given a CFT2, a $3d$ gravity is induced, in which a $3d$ QFT lives. This $3d$ QFT in turn induces a $4d$ gravity, and so on.
We showed that the metric $g_{ij}=-[\chi_{ij}]$ is a derived but not a fundamental quantity. Perhaps this is why gravity is non-renormalizable and we should not try to quantize the metric itself.
We addressed the vacuum Einstein’s equation in this work. From the derivations, including matter sources is a straightforward generalization.
In physics, energy dominates and links everything. But now, we tend to believe the fundamental object in physics is entropy.
Acknowledgements.
Acknowledgments.
This work is supported in part by NSFC (Grant No. 12105191, 12275183 and 12275184).